Operations Report – April 19th

Crew 280 Operations Report 19-04-2023

SOL: 3

Name of person filing report: Cesca Cufí (Crew Engineer) and Neus Sabaté (Mission Specialist)

Non-nominal systems:

Robotic observatory. Notes: replacement parts are scheduled to be installed during crew 261 rotation.

Rover Spirit. Notes: corroded battery wire. Replacement pending (Sergii). Rover can only be used for short distance EVA.


Spirit rover used: no

Hours: 216.7

Beginning charge: –

Ending charge: –

Currently charging: no

Opportunity rover used: no

Hours: 120.7

Beginning charge: –

Ending charge: –

Currently charging: no

Curiosity rover used: no

Hours: 229.8

Beginning charge: –

Ending charge: –

Currently charging: no

Perseverance rover used: no

Hours: 261.1

Beginning charge: –

Ending charge: –

Currently charging: no

General notes on rovers: none

Summary of Hab operations:

Water use: 20 gal (from 9AM Sol2 to 8AM Sol3)

Water (static tank): 443 gal (at 6:30PM). Water tank refilled this afternoon.

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: used from 18:00 to 21:00 for personal communication and reports.

Summary of suits and radios:

Suits: not used.

Radios: EVA radios not used.Station radios

All radio and spacesuit batteries checked before EVA: ok.

Summary of GreenHab operations:

Water use: 1,5 gal for experiments in the Science Dome

Heater: off

Supplemental light: off

Harvest: 3g of mint and 263g of tomatoes

Summary of Science Dome operations:

Laia Ribas (Green Hab Officer) continued preparing her experiments.

Neus Sabaté (Mission Specialist) started her experiment.

Dual split: off

Summary of RAM operations:

Cesca Cufí (Crew Engineer) continued her project integration.

Summary of any observatory issues: none

Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.

Summary of health and safety issues: we have replaced batteries of portable gas detectors. The exhausted batteries have been placed in a ziplog bag and left next to the B and NB bins in the lower deck.

Summary of other system’s issues:

Hab heater: it turned off about 2PM. No reason found. Crew Engineer and Commander troubleshoot the issue with the help of mission control. The heater showed the following Light code message: one short blink followed by 3 long blinks. The following actions were executed following mission control orders. The heater was turned off and on again, without results. Then the heater switch on the voltage panel was turned off and on again. The heater LED indicator was then constantly on, meaning that the heater was receiving 24VAC. After a few seconds, the heater turned on.

At 6.45PM the heater turned off again and it was troubleshot the same way.

Some bridles of the corridor tents were missing or broken after overnight winds. We replaced the damage with new bridles.

Questions, concerns and requests to Mission Support:

Spirit rover: replacement of then corroded wire

Kitchen: knobs replacement

Science dome: we have been organizing the science dome in order to fit all the planned experiments of the mission. Yesterday we had to move the microwave from one of the tables in order to have more dedicated space for biology experiments. Today, we allocated it in its original place but we realized that that space was needed to run the genome sequencing experiment. Therefore, we request to Mission Support the permission to store the microwave in a safe place during the mission and put it back to its original location once the experiments are completed. Attached you will find a pdf that explains how the different available locations of the science dome will be used and the suggested temporary location for the microwave.

Copyright © The Mars Society. All rights reserved. | Main Site